how do I rewrite the quadratic equation x^2+6x-8 in vertex form?
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



What equation? Your quadratic trinomial isn't equal to anything, so you don't have an equation. Quadratic equation is the wrong terminology anyway. You want to talk about a quadratic function if you want to convert to vertex form.

Try this:



Now, complete the square on the right and then collect terms to result in a function that has the following pattern:



which is a parabola with a vertical axis of symmetry and a vertex of (h,k). The parabola opens upward if , downward otherwise. is the distance from the vertex to the focus of the parabola and is the distance from the vertex to the directrix.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

how do I rewrite the quadratic equation x^2+6x-8 in vertex form?
-------
Complete the square:
x^2+6x-8 = y
x^2+6x+3^2 = y+8+3^2
----
(x+3)^2 = y+17
----------------------
Vertex: (-3,-17)
=======================
Cheers,
Stan H.
=======================